# Categorical Time-Reversal Symmetries

### 30 Mar 2026 v1

@@ V.6.4 ${\mathbb{Z}}_{4}^{T}\to{\mathbb{Z}}_{2}$ partial SSB
We see that this category can be renamed as $\mathsf{Vec}_{{\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}^{T}}^{\alpha}$ with $\alpha$ given as above. Notice that this cocyle $\alpha$ is a nontrivial one since by inspection it is exactly the generator of
@@ V.7 $S_{3}^{T}={\mathbb{Z}}_{3}\rtimes{\mathbb{Z}}_{2}^{T}$
  • $S_{3}^{T}\to{\mathbb{Z}}_{2}^{T}$ partial SSB. Since $H^{2}({\mathbb{Z}}_{2}^{T},U(1))={\mathbb{Z}}_{2}$, we can further put a Haldane chain onto the vacuua. The two phases are denoted as $({\mathbb{Z}}_{2}^{T},1),\penalty 10000\ ({\mathbb{Z}}_{2}^{T},\omega)$.
@@ V.8 Gauging Finite Subgroups
Compare with the result of gauging central subgroup for a unitary group, the main new feature here is the semi-direct product $\widehat{N}\rtimes K^{T}$. The anomaly $\omega\in H^{3}(G^{T}/N\ltimes\widehat{N},U(1)_{T})$ lives in the subgroup
@@
Notice since $G^{T}/N$ acts both on $\widehat{N}$ and the coefficient $U(1)_{T}$ as inversion, its action on $H^{1}(\widehat{N},U(1))\simeq N$ is trivial. Then $\omega$ can be viewed as the extension class $e_{2}\in H^{2}(G^{T}/N,N)$ itself. Hence $\omega$ is nontrivial if and only if $N\to G^{T}\to G^{T}/N$ is a non-split extension.
@@ VI.1 T-Enrichment versus Gauging
For a (d+1)(d+1)D internal finite $G$-symmetry, the standard SymTFT paradigm identifies the SymTFT as a (d+2)(d+2)D $G$-gauge theory. Following this principle one tries to define the SymTFT for an anti-linear symmetry, e.g. the time-reversal ${\mathbb{Z}}_{2}^{T}$, as the (d+2)(d+2)D ${\mathbb{Z}}_{2}^{T}$-gauge theory”. However there is a fundamental obstruction for gauging time-reversal. Coupling a field theory to time-reversal background is really putting the theory on an unoriented manifold $M$, whose first Stiefel–Whitney class $w_{1}(M)$ can be thought of as the time-reversal background field. However, unlike the background gauge field for an internal symmetry, the time-reversal background field $w_{1}(M)$ is determined by the underlying manifold $M$ and we are not allowed to vary it. Hence the only meaningful way of dynamically gauging time-reversal is to sum over spacetime topologies, which forces us into the realm of quantum gravity. Since we would like to stay in the realm of (topological) quantum field theory, we will not take this approach for the SymTFT for ${\mathbb{Z}}_{2}^{T}$, or any anti-linear symmetry. Instead, we argue that it suffices to take an unoriented TFT, or in condensed matter language, a ${\mathbb{Z}}_{2}^{T}$-enriched topological order, as our SymTFT for an anti-linear symmetry. We will show that for a ${\mathbb{Z}}_{2}^{T}$-enriched non-chiral topological order, its ${\mathbb{Z}}_{2}^{T}$-SSB boundaries are described by Galois-real fusion categories, hence give rise to anti-linear symmetry categories.
@@ VI.2.2 The Theory of Non-chiral 3d ${\mathbb{Z}}_{2}^{T}$-SETs
A subtlety here is that without time-reversal enrichment (or any symmetry), the Drinfeld center description $\mathcal{Z}(\mathcal{C})$ for a non-chiral topological is equivalent to the bimodule 2-category description $\mathsf{Bimod}(\mathcal{C})$, because the latter is simply $\Sigma\mathcal{Z}(\mathcal{C})$. However with time-reversal enrichment and work over $\mathbb{R}$, this is no longer true in general. Namely there could be codimension-1 defects in the bulk of a non-chiral ${\mathbb{Z}}_{2}^{T}$-SET that does not arise as a condensation defect of codimension-2 defects. This happens when the ${\mathbb{Z}}_{2}^{T}$-SET is a ${\mathbb{Z}}_{2}^{T}$-SSB order. As long as the input real fusion category $\mathcal{C}$ is not a complex fusion category(i.e., either it is $\mathbb{R}$-real, or Galois-real with nonzero $\mathcal{C}_{T}$ component), then $\mathsf{Bimod}_{\mathbb{R}}(\mathcal{C})=\Sigma\mathcal{Z}(\mathcal{C})$ remains true  [^5] Proof of this will appear elsewhere. and there is no harm in saying “the bulk is completely determined by the Drinfeld center”. Therefore in what follows we consider two types of input real fusion categories for a ${\mathbb{Z}}_{2}^{T}$-SET: either an $\mathbb{R}$-real fusion category, or a Galois-real fusion category $\mathcal{C}_{1}\oplus\mathcal{C}_{T}$ with nonzero $\mathcal{C}_{T}$. We will then simply say the Drinfeld center $\mathcal{Z}(\mathcal{C})$ is the ${\mathbb{Z}}_{2}^{T}$-SET.
Given a real fusion category $\mathcal{C}\in 3\mathsf{Vec}_{\mathbb{R}}$, a ${\mathbb{Z}}_{2}^{T}$-symmetric boundary condition for the ${\mathbb{Z}}_{2}^{T}$-SET defined by $\mathcal{C}$ is simply a morphism in $\text{Hom}_{3\mathsf{Vec}_{\mathbb{R}}}(\mathsf{Vec},\mathcal{C})$, which is a module category over $\mathcal{C}$. Therefore the ${\mathbb{Z}}_{2}^{T}$-symmetric boundaries  [^6] By a ${\mathbb{Z}}_{2}^{T}$-symmetric boundary we mean one that does not break the enriching ${\mathbb{Z}}_{2}^{T}$ explicitly, but we allow for spontaneously breaking ${\mathbb{Z}}_{2}^{T}$. are classified by module categories over the input real fusion category. In particular, for every input real fusion category $\mathcal{C}$, there is a canonical (Dirichlet) boundary given by the regular module $\mathcal{C}$ itself, and the real fusion category of ${\mathbb{Z}}_{2}^{T}$-symmetric defects on the canonical boundary is
@@
This implies the underlying gapped boundary is in fact multi-fusion with $\mathbf{1}=\mathbf{1}_{+}\oplus\mathbf{1}_{-}$. The two ground states are permuted by ${\mathbb{Z}}_{2}^{T}$-action, hence there is still only one ${\mathbb{Z}}_{2}^{T}$-symmetric ground state. The Galois-real fusion category $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{T}$ on the boundary should be viewed as the category of either defects within each vacuum, or domain walls between the two vacuua.
@@ VI.3 Alternative Description of ${\mathbb{Z}}_{2}^{T}$-SETs
Figure 5: T-enriched SymTFT quiche for the Galois-real category $\mathcal{C}_{1}\oplus\mathcal{C}_{T}$: The fusion category $\mathcal{C}_{1}$ ($\mathcal{C}_{1}^{*}$) is the canonical boundary of the Levin-Wen topological order $\mathfrak{Z}(\mathcal{C}_{1})$ ($\mathfrak{Z}(\mathcal{C}_{1}^{*})$). The ${\mathbb{Z}}_{2}^{T}$ (anti-unitary) automorphism defect in the bulk of the SymTFT ends in $\mathcal{C}_{T}$, which is an invertible $\mathcal{C}_{1}$-$\mathcal{C}_{1}^{*}$ bimodule. The boundary of the $D_{T}$ surface is a domain wall between two vacuua on the boundary (denoted $|\uparrow\rangle$ and $|\downarrow\rangle$).
@@ VII.1 T-Enriched SymTFT Quiches
The surface $D_{T}$ then ends on the symmetry boundary as a nontrivial defect between $\mathcal{C}_{1}$ and $\mathcal{C}_{1}^{*}$, with potentially multiple boundary conditions. It is always understood that the boundary has two vacuua permuted by the $D_{T}$-action. If $\mathcal{C}_{1}$ is the category of topological defects in one vacuua, while $\mathcal{C}_{1}^{*}$ is the category of topological defects in the other vacuua. The collection of all boundary conditions for the $D_{T}$ surface form a $\mathcal{C}_{1}-\mathcal{C}_{1}^{*}$ bimodule category $\mathcal{C}_{T}$. Fix any boundary condition $T\in\mathcal{C}_{T}$, other boundary conditions are always obtained by fusing with orientation-preserving defects in $\mathcal{C}_{1}$ or $\mathcal{C}_{1}^{*}$. We will denote by $a\rhd x,\penalty 10000\ a\in\mathcal{C}_{1},\penalty 10000\ x\in\mathcal{C}_{T}$ fusing with $a$ from above , and $x\lhd a$ for fusing with $a$ from below. A defect in $\mathcal{C}_{1}$(or $\mathcal{C}_{1}^{*}$) will be called a linear defect, and a domain wall between the two vacuua on the boundary(where $D_{T}$ ends) is called an anti-linear defect.
@@ VII.2 T-Enriched Quiche for ${\mathbb{Z}}_{2}^{T}$
The boundary $T$ of the bulk $D_{T}$ surface is a domain wall between the two (trivial) vacuua on the boundary.
@@ VII.5 T-Enriched Quiche for Galois-real TY-categories
which is nothing but the condition that $q$ viewed as a bicharacter of $A$ is skew-symmetric. We conclude that a ${\mathbb{Z}}_{2}^{T}$-enrichment on $\mathcal{Z}(\mathsf{Vec}_{A})$ that gives rise to a Galois-real TY symmetry category on the $\mathsf{Vec}_{A}$-boundary is determined by a non-degenerate skew-symmetric bicharacter $q$ of $A$. Note that for the usual complex TY categories, the Frobenius-Schur indicator can be viewed as the freedom of stacking with a (2+1)(2+1)d ${\mathbb{Z}}_{2}$-SPT in the bulk before gauging a chare-flux exchange ${\mathbb{Z}}_{2}$-symmetry. However here $H^{3}({\mathbb{Z}}_{2}^{T},U(1)_{T})=0$ hence there is no more freedom. This explains the absence of a Frobenius-Schur indicator in the classification of Galois-real TY categories.
@@ VIII Conclusions and Outlook
We have put forward a theory for categorical symmetries including time-reversal action ${\mathbb{Z}}_{2}^{T}$, and made a case that the natural mathematical habitat for such symmetries are real fusion categories. In fact, symmetry categories with time-reversal have to be so-called Galois-real fusion categories, as opposed to $\mathbb{R}$-real ones (where the endomorphisms of the identity form a real vector space). Generalized gauging (or Morita equivalence between categories) in the world of real fusion categories does not always follow the intuition that we may have formed working with complex fusion categories, and we provided some key examples to illustrate this. E.g. the symmetries $A\times{\mathbb{Z}}_{2}^{T}$ and $A\rtimes{\mathbb{Z}}_{2}^{T}$ are gauge related and even the representation category of an abelian group ${\mathbb{Z}}_{n}^{T}$ can be non-invertible. We furthermore set up a $T$-enriched SymTFT framework for such real fusion categories, which in particular gives an elegant way to derive many of the Morita equivalence relations.
Although the general theory we put forward is applicable in any dimensions, our examples focus on (1+1)(1+1)d real fusion categories, and finite many simple objects. It would be interesting to extend our analysis to a full study of the phases using the SymTFT and to symmetries in higher dimensions, including continuous symmetries, such as ${\mathbb{Z}}_{4}^{T},U(1)\rtimes{\mathbb{Z}}_{2}^{T},U(1)\times{\mathbb{Z}}_{2}^{T}$ that play a role in topological insulator/superconductors. One challenge for developing the SymTFT sandwich construction for gapped phases is the ${\mathbb{Z}}_{2}^{T}$-symmetry breaking on the symmetry boundary. This could be avoided by considering instead ${\mathbb{Z}}_{2}^{T}$-symmetric (non-SSB) boundaries as in [19], which may lead to a complete sandwich construction for gapped phases with symmetries of the form $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$, up to pure ${\mathbb{Z}}_{2}^{T}$-SPTs. Alternatively, one could interpret the compactification of the sandwich with a ${\mathbb{Z}}_{2}^{T}$-SSB boundary not as the gapped phase itself but a doubled version where the two vacuua are interpreted as the system itself and its complex conjugate.

### 02 Apr 2026 v2

@@ V.6.4 ${\mathbb{Z}}_{4}^{T}\to{\mathbb{Z}}_{2}$ partial SSB
We see that this category can be renamed as $\mathsf{Vec}_{{\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}^{T}}^{\alpha}$ with $\alpha$ given as above. Notice that this cocycle $\alpha$ is a nontrivial one since by inspection it is exactly the generator of
@@ V.7 $S_{3}^{T}={\mathbb{Z}}_{3}\rtimes{\mathbb{Z}}_{2}^{T}$
  • $S_{3}^{T}\to{\mathbb{Z}}_{2}^{T}$ partial SSB. Since $H^{2}({\mathbb{Z}}_{2}^{T},U(1))={\mathbb{Z}}_{2}$, we can further put a Haldane chain onto the vacua. The two phases are denoted as $({\mathbb{Z}}_{2}^{T},1),\penalty 10000\ ({\mathbb{Z}}_{2}^{T},\omega)$.
@@ V.8 Gauging Finite Subgroups
Compared with the result of gauging central subgroup for a unitary group, the main new feature here is the semi-direct product $\widehat{N}\rtimes K^{T}$. The anomaly $\omega\in H^{3}(G^{T}/N\ltimes\widehat{N},U(1)_{T})$ lives in the subgroup
@@
Notice that, since $G^{T}/N$ acts both on $\widehat{N}$ and the coefficient $U(1)_{T}$ as inversion, its action on $H^{1}(\widehat{N},U(1))\simeq N$ is trivial. Then $\omega$ can be viewed as the extension class $e_{2}\in H^{2}(G^{T}/N,N)$ itself. Hence $\omega$ is nontrivial if and only if $N\to G^{T}\to G^{T}/N$ is a non-split extension.
@@ VI.1 T-Enrichment versus Gauging
For a (d+1)(d+1)D internal finite $G$-symmetry, the standard SymTFT paradigm identifies the SymTFT as a (d+2)(d+2)D $G$-gauge theory. Following this principle one tries to define the SymTFT for an anti-linear symmetry, e.g. the time-reversal ${\mathbb{Z}}_{2}^{T}$, as the (d+2)(d+2)D ${\mathbb{Z}}_{2}^{T}$-gauge theory”. However, there is a fundamental obstruction for gauging time-reversal. Coupling a field theory to time-reversal background is really putting the theory on an unoriented manifold $M$, whose first Stiefel–Whitney class $w_{1}(M)$ can be thought of as the time-reversal background field. However, unlike the background gauge field for an internal symmetry, the time-reversal background field $w_{1}(M)$ is determined by the underlying manifold $M$ and we are not allowed to vary it. Hence the only meaningful way of dynamically gauging time-reversal is to sum over spacetime topologies, which forces us into the realm of quantum gravity. Since we would like to stay in the realm of (topological) quantum field theory, we will not take this approach for the SymTFT for ${\mathbb{Z}}_{2}^{T}$, or any anti-linear symmetry. Instead, we argue that it suffices to take an unoriented TFT, or in condensed matter language, a ${\mathbb{Z}}_{2}^{T}$-enriched topological order, as our SymTFT for an anti-linear symmetry. We will show that for a ${\mathbb{Z}}_{2}^{T}$-enriched non-chiral topological order, its ${\mathbb{Z}}_{2}^{T}$-SSB boundaries are described by Galois-real fusion categories, hence give rise to anti-linear symmetry categories.
@@ VI.2.2 The Theory of Non-chiral 3d ${\mathbb{Z}}_{2}^{T}$-SETs
A subtlety here is that without time-reversal enrichment (or any symmetry), the Drinfeld center description $\mathcal{Z}(\mathcal{C})$ for a non-chiral topological is equivalent to the bimodule 2-category description $\mathsf{Bimod}(\mathcal{C})$, because the latter is simply $\Sigma\mathcal{Z}(\mathcal{C})$. However with time-reversal enrichment and work over $\mathbb{R}$, this is no longer true in general. Namely there could be codimension-1 defects in the bulk of a non-chiral ${\mathbb{Z}}_{2}^{T}$-SET that does not arise as a condensation defect of codimension-2 defects. This happens when the ${\mathbb{Z}}_{2}^{T}$-SET is a ${\mathbb{Z}}_{2}^{T}$-SSB order  [^5] For instance, take $\mathcal{C}=\mathsf{Vec}_{\mathbb{C}}$, then $\mathcal{Z}(\mathsf{Vec}_{\mathbb{C}})=\mathsf{Vec}_{\mathbb{C}}$ while $\mathsf{Bimod}_{\mathbb{R}}(\mathsf{Vec}_{\mathbb{C}})=\text{End}_{\mathbb{R}}(2\mathsf{Vec}_{\mathbb{C}})=2\mathsf{Vec}_{\mathbb{C}}\oplus 2\mathsf{Vec}_{\mathbb{C}}$ is not connected. This is the trivial (2+1)(2+1)d ${\mathbb{Z}}_{2}^{T}$-SSB order. As long as the input real fusion category $\mathcal{C}$ is not a complex fusion category(i.e., either it is $\mathbb{R}$-real, or Galois-real with nonzero $\mathcal{C}_{T}$ component), then $\mathsf{Bimod}_{\mathbb{R}}(\mathcal{C})=\Sigma\mathcal{Z}(\mathcal{C})$ remains true  [^6] Proof of this will appear elsewhere. and there is no harm in saying “the bulk is completely determined by the Drinfeld center”. Therefore in what follows we consider two types of input real fusion categories for a ${\mathbb{Z}}_{2}^{T}$-SET: either an $\mathbb{R}$-real fusion category, or a Galois-real fusion category $\mathcal{C}_{1}\oplus\mathcal{C}_{T}$ with nonzero $\mathcal{C}_{T}$. We will then simply say the Drinfeld center $\mathcal{Z}(\mathcal{C})$ is the ${\mathbb{Z}}_{2}^{T}$-SET.
Given a real fusion category $\mathcal{C}\in 3\mathsf{Vec}_{\mathbb{R}}$, a ${\mathbb{Z}}_{2}^{T}$-symmetric boundary condition for the ${\mathbb{Z}}_{2}^{T}$-SET defined by $\mathcal{C}$ is simply a morphism in $\text{Hom}_{3\mathsf{Vec}_{\mathbb{R}}}(\mathsf{Vec}_{\mathbb{R}},\mathcal{C})$, which is a module category over $\mathcal{C}$. Therefore the ${\mathbb{Z}}_{2}^{T}$-symmetric boundaries  [^7] By a ${\mathbb{Z}}_{2}^{T}$-symmetric boundary we mean one that does not break the enriching ${\mathbb{Z}}_{2}^{T}$ explicitly, but we allow for spontaneously breaking ${\mathbb{Z}}_{2}^{T}$. are classified by module categories over the input real fusion category. In particular, for every input real fusion category $\mathcal{C}$, there is a canonical (Dirichlet) boundary given by the regular module $\mathcal{C}$ itself, and the real fusion category of ${\mathbb{Z}}_{2}^{T}$-symmetric defects on the canonical boundary is
@@
This implies the underlying gapped boundary is in fact multi-fusion with $\mathbf{1}=\mathbf{1}_{+}\oplus\mathbf{1}_{-}$. The two ground states are permuted by ${\mathbb{Z}}_{2}^{T}$-action, hence there is still only one ${\mathbb{Z}}_{2}^{T}$-symmetric ground state. The Galois-real fusion category $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{T}$ on the boundary should be viewed as the category of either defects within each vacuum, or domain walls between the two vacua.
@@ VI.3 Alternative Description of ${\mathbb{Z}}_{2}^{T}$-SETs
Figure 5: T-enriched SymTFT quiche for the Galois-real category $\mathcal{C}_{1}\oplus\mathcal{C}_{T}$: The fusion category $\mathcal{C}_{1}$ ($\mathcal{C}_{1}^{*}$) is the canonical boundary of the Levin-Wen topological order $\mathfrak{Z}(\mathcal{C}_{1})$ ($\mathfrak{Z}(\mathcal{C}_{1}^{*})$). The ${\mathbb{Z}}_{2}^{T}$ (anti-unitary) automorphism defect in the bulk of the SymTFT ends in $\mathcal{C}_{T}$, which is an invertible $\mathcal{C}_{1}$-$\mathcal{C}_{1}^{*}$ bimodule. The boundary of the $D_{T}$ surface is a domain wall between two vacua on the boundary (denoted $|\uparrow\rangle$ and $|\downarrow\rangle$).
@@ VII.1 T-Enriched SymTFT Quiches
The surface $D_{T}$ then ends on the symmetry boundary as a nontrivial defect between $\mathcal{C}_{1}$ and $\mathcal{C}_{1}^{*}$, with potentially multiple boundary conditions. It is always understood that the boundary has two vacua permuted by the $D_{T}$-action. If $\mathcal{C}_{1}$ is the category of topological defects in one vacua, while $\mathcal{C}_{1}^{*}$ is the category of topological defects in the other vacua. The collection of all boundary conditions for the $D_{T}$ surface form a $\mathcal{C}_{1}-\mathcal{C}_{1}^{*}$ bimodule category $\mathcal{C}_{T}$. Fix any boundary condition $T\in\mathcal{C}_{T}$, other boundary conditions are always obtained by fusing with orientation-preserving defects in $\mathcal{C}_{1}$ or $\mathcal{C}_{1}^{*}$. We will denote by $a\rhd x,\penalty 10000\ a\in\mathcal{C}_{1},\penalty 10000\ x\in\mathcal{C}_{T}$ fusing with $a$ from above , and $x\lhd a$ for fusing with $a$ from below. A defect in $\mathcal{C}_{1}$(or $\mathcal{C}_{1}^{*}$) will be called a linear defect, and a domain wall between the two vacua on the boundary(where $D_{T}$ ends) is called an anti-linear defect.
@@ VII.2 T-Enriched Quiche for ${\mathbb{Z}}_{2}^{T}$
The boundary $T$ of the bulk $D_{T}$ surface is a domain wall between the two (trivial) vacua on the boundary.
@@ VII.5 T-Enriched Quiche for Galois-real TY-categories
which is nothing but the condition that $q$ viewed as a bicharacter of $A$ is skew-symmetric. We conclude that a ${\mathbb{Z}}_{2}^{T}$-enrichment on $\mathcal{Z}(\mathsf{Vec}_{A})$ that gives rise to a Galois-real TY symmetry category on the $\mathsf{Vec}_{A}$-boundary is determined by a non-degenerate skew-symmetric bicharacter $q$ of $A$. Note that for the usual complex TY categories, the Frobenius-Schur indicator can be viewed as the freedom of stacking with a (2+1)(2+1)d ${\mathbb{Z}}_{2}$-SPT in the bulk before gauging a charge-flux exchange ${\mathbb{Z}}_{2}$-symmetry. However, here $H^{3}({\mathbb{Z}}_{2}^{T},U(1)_{T})=0$ hence there is no more freedom. This explains the absence of a Frobenius-Schur indicator in the classification of Galois-real TY categories.
@@ VIII Conclusions and Outlook
We have put forward a theory for categorical symmetries including time-reversal action ${\mathbb{Z}}_{2}^{T}$, and made a case that the natural mathematical habitat for such symmetries is real fusion categories. In fact, symmetry categories with time-reversal have to be so-called Galois-real fusion categories, as opposed to $\mathbb{R}$-real ones (where the endomorphisms of the identity form a real vector space). Generalized gauging (or Morita equivalence between categories) in the world of real fusion categories does not always follow the intuition that we may have formed working with complex fusion categories, and we provided some key examples to illustrate this. E.g. the symmetries $A\times{\mathbb{Z}}_{2}^{T}$ and $A\rtimes{\mathbb{Z}}_{2}^{T}$ are gauge related and even the representation category of an abelian group ${\mathbb{Z}}_{n}^{T}$ can be non-invertible. We furthermore set up a $T$-enriched SymTFT framework for such real fusion categories, which in particular gives an elegant way to derive many of the Morita equivalence relations.
Although the general theory we put forward is applicable in any dimensions, our examples focus on (1+1)(1+1)d real fusion categories, and finite many simple objects. It would be interesting to extend our analysis to a full study of the phases using the SymTFT and to symmetries in higher dimensions, including continuous symmetries, such as ${\mathbb{Z}}_{4}^{T},U(1)\rtimes{\mathbb{Z}}_{2}^{T},U(1)\times{\mathbb{Z}}_{2}^{T}$ that play a role in topological insulator/superconductors. One challenge for developing the SymTFT sandwich construction for gapped phases is the ${\mathbb{Z}}_{2}^{T}$-symmetry breaking on the symmetry boundary. This could be avoided by considering instead ${\mathbb{Z}}_{2}^{T}$-symmetric (non-SSB) boundaries as in [19], which may lead to a complete sandwich construction for gapped phases with symmetries of the form $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$, up to pure ${\mathbb{Z}}_{2}^{T}$-SPTs. Alternatively, one could interpret the compactification of the sandwich with a ${\mathbb{Z}}_{2}^{T}$-SSB boundary not as the gapped phase itself but a doubled version where the two vacua are interpreted as the system itself and its complex conjugate.